# Diff-DAgger: Uncertainty Estimation with Diffusion Policy for Robotic Manipulation

### v2

@@ Abstract
Recently, diffusion policy has shown impressive results in handling multi-modal tasks in robotic manipulation. However, it has fundamental limitations in out-of-distribution failures that persist due to compounding errors and its limited capability to extrapolate. One way to address these limitations is robot-gated DAgger, an interactive imitation learning with a robot query system to actively seek expert help during policy rollout. While robot-gated DAgger has high potential for learning at scale, existing methods like Ensemble-DAgger struggle with highly expressive policies: They often misinterpret policy disagreements as uncertainty at multi-modal decision points. To address this problem, we introduce Diff-DAgger, an efficient robot-gated DAgger algorithm that leverages the training objective of diffusion policy. We evaluate Diff-DAgger across different robot tasks including stacking, pushing, and plugging, and show that Diff-DAgger improves the task failure prediction by 37%, the task completion rate by 14%, and reduces the wall-clock time by up to 540%. We hope that this work opens up a path for efficiently incorporating expressive yet data-hungry policies into interactive robot learning settings. Project website: diffdagger.github.io
@@ I Introduction
However, none of the existing robot-gated DAgger approaches fully addresses the multi-modality present in demonstration data. For example, the most widely adopted Ensemble-DAgger approaches use the action variance among an ensemble of policies to estimate uncertainty [6, 7, 8]. This means that when tasks admit multiple solutions, policies might erroneously indicate high uncertainty even when in a previously visited state. We illustrate this problem with a 2D navigation example (Figure 2 (a)), where two primary paths—clockwise and counterclockwise around a circle—lead to the goal. This scenario allows us to categorize regions as in-distribution uni-modal (along the circular paths, shown as mixed red and orange), in-distribution multi-modal (at the path divergence point, shown as green), and out-of-distribution (significant deviations from expected paths, shown as blue). While ensemble methods for robot-gated DAgger effectively distinguish in-distribution uni-modal from out-of-distribution data using action variance, they fail at in-distribution multi-modal points due to policy disagreement (Figure 2 (b)).
@@
Figure 2: (a) Visualization of 2D navigation toy example. The orange dots represent the mode for following a clockwise path around the circle and the red dots represent another mode for following a counter-clockwise path. The green dots represent the multi-modal region where multiple diverging paths exist. The blue region indicates out-of-distribution. (b) Comparison of action variance and diffusion loss for different regions across 100 runs in the 2D navigation example. The left bar plot displays the standard deviation of action across policies for in-distribution uni-modal (ID, UM), in-distribution multi-modal (ID, MM), and out-of-distribution (OOD) regions. The right bar plot shows the corresponding diffusion loss for these regions.
Incorporating an expressive policy class that handles multi-modal data into the DAgger framework is a promising but under-explored direction. In light of this, we propose Diff-DAgger, an efficient robot-gated DAgger algorithm that addresses multi-modality in demonstration data while reducing wall-clock time compared to other methods. Diff-DAgger uses a query system based on the loss function of the diffusion model, referred to as diffusion loss, which significantly improves the effectiveness of robot queries across various evaluation metrics. In three challenging manipulation tasks, Diff-DAgger achieves a 37% higher F1 score in predicting task failure, and the collected data leads to 14% improvements in policy performance while improving the wall clock time by 540% in comparison to the baseline methods.
We make the following contributions: First, we develop a novel robot-gated query system based on diffusion loss to handle multi-modality in expert demonstration. Second, we quantitatively evaluate Diff-DAgger’s efficiency in terms of task failure prediction, task completion, and wall-clock time during interactive learning to demonstrate its superiority over other baseline methods. Lastly, we discuss design choices that accelerate policy learning to enable faster convergence and improved overall performance.
@@ II-B Uncertainty Estimation of Robot Action
The most widely used method for estimating action uncertainty is to use an ensemble of policies bootstrapped on a subset of the training data and using the level of disagreement among the policies output to estimate uncertainty [6, 7, 8]. However, the ensemble method for uncertainty estimation can be problematic in practice, especially when the task allows multiple modes of completion. Specifically, they are prone to either mistaking in-distribution multi-modal data as out-of-distribution data or vice versa. This is because expressive ensemble policies may disagree due to the presence of multiple viable strategies where each policy is certain about its generated plan, leading to a high action variance value. Furthermore, if the policy generates an action sequence instead of a single action, the uncertainty estimation may explode due to increasing divergence over time or lose accuracy if a shorter horizon is used for calculating action variance.
@@
In contrast to the existing robot-gated DAgger methods, our approach addresses multi-modality in the demonstration data by incorporating the loss function of diffusion policy into the decision rule. This approach does not need multiple policies for uncertainty estimation, reducing both training and inference time while benefiting from using an expressive policy class.
@@ III-B1 Action Diffusion
In diffusion policy, robot behavior is generated through an iterative denoising process based on Langevin dynamics, formally modeled as a Denoising Diffusion Probabilistic Model (DDPM) [2]. This denoising process begins with a Gaussian prior, $\mathcal{N}(0,I)$, and iteratively removes noise to reconstruct the original data. Diffusion policy conditions on observation sequences and generates an action sequence, achieving state-of-the-art performance across various robotic tasks [3].
@@ III-B2 Loss Function
During training, diffusion policy samples a denoising timestep and a Gaussian noise vector to corrupt the action sequence. It then predicts the target objective given the corrupted action sequence and the observation sequence, and performs optimization by minimizing the following loss function:
$L_{\pi}(o,a,\epsilon,t)=\|f(\epsilon,a)-f_{\theta}(o,\sqrt{\alpha_{t}}a+\sqrt{1-\alpha_{t}}\epsilon,t)\|^{2}$ (1)
where $o$ is the observation sequence, $a$ is the action sequence, and $\epsilon$ is a noise vector from Gaussian distribution $\mathcal{N}(0,I)$. The parameter $\alpha_{t}$ is a time-dependent factor that modulates the influence of the original data and the added noise during the diffusion process, and $t$ is the denoising timestep. The loss function $L(o,a,\epsilon,t)$ calculates the mean squared error (MSE) between the target variable $f(\epsilon,a)$, which is a function of noise and ground truth action depending on the choice of target prediction, and the noise predicted by the model $f_{\theta}$, given the noised data at time $t$ [18].
@@ III-C1 OOD Detection via Diffusion Loss
15:         if CDF(loss) $<\alpha$ for last $K$ timesteps then
@@ III-C2 Data Aggregation and Policy Update
We now summarize the data aggregation and policy update procedures. Diff-DAgger first initializes the policy $\pi_{r}$ by training it on Ni initial expert demonstrations. After training the policy, the robot continues rolling out the action in the environment until high diffusion loss is detected. When the robot query is made, experts intervene from the current scene and complete the task. Upon completion of the task by either the robot or the expert, the task is initialized to a random initial configuration. If the robot neither queries nor completes the task after a time limit, automatic expert intervention takes place. After every Nd expert intervention has been made, we perform a policy update, which consists of policy training and threshold setting. The policy training involves training the policy from scratch, and threshold setting involves setting the base values from the training dataset for the OOD violation threshold used during robot-query decision-making. The full pseudo-code is available in Algorithm 1.
@@ IV-A Experiment Setup
Figure 3: Tasks (Bottom) and Camera Views (Top). For pushing, a single camera is used. For stacking and plugging, three cameras are used to account for occlusion and gripper motion.
Task # Obj Contact Rich High-Fidel TL Expert Type Task Dur Succ Rate
Stacking 2 False False 250 Motion Planner 60.3 1.00
Pushing 1 True False 250 RL Agent 65.7 0.93
Plugging 2 False True 500 Motion Planner 147 0.98
TABLE I: Task and Expert Summary. # Obj: number of objects, Contact-rich: contact-richness between robot and objects, High-Fidel: High-fidelity task, TL: time limit in simulation steps at 20 Hz, Expert Type: type of expert, Task Dur: average expert task duration in simulation steps, Exp SR: expert success rate during intervention.
@@ IV-A1 Tasks and Experts
We include three tasks in simulation—stacking, pushing, and plugging—using the ManiSkill environment [19]. In the stacking task, the robot must pick up a red cube and place it on top of a green cube, testing its perception and basic manipulation ability. The pushing task requires the robot to push a red tee into a predefined position using an end-effector stick, presenting a challenge of contact-richness and explicit multi-modality in task completion. Finally, the plugging task requires picking up a charger and plugging it into a receptacle, posing a significant challenge of fine-manipulation with with millimeter-level precision.
We design experts as either multi-modal or long-horizon planners to evaluate different expert types. Specifically, we use RL agents and motion planners as the experts: the motion planner, is a long-horizon planner that generates actions based on future goals, whereas multiple RL agents with different reward functions provide different modes for task completion. For stacking and plugging tasks, we use a rule-based RRT motion planner to simulate the expert [20]. For pushing, two experts are trained using PPO algorithm [21] with varying rewards: in addition to task-related reward, one agent is additionally rewarded for rotating the object clockwise direction, whereas other agent is rewarded for rotating counter-clockwise (2 Experts). We additionally compare learning from the counter-clockwise agent only (1 Expert). The tasks and experts are summarized in Table I.
@@ IV-A2 Baselines
IV-A3 Model Details
In all tasks, we use the CNN-based [22] diffusion policy with FiLM conditioning [23] for the observation. In our study, we test both state-based and image-based data: For the state-based case, we use the internal robot proprioception data (robot joint position and end-effector pose) and the ground-truth pose of the external objects as the input to the model. For the image-based case, we use a vision encoder to extract the embedding vector from the image and concatenate it with the proprioception data. Additionally, we use spatial-softmax instead of average pooling for the vision encoder [24]. All methods use position-based control, either with joint position control or end-effector pose control. For end-effector pose control, 6D representation for the orientation is used [25].
Task ND Method TP TN FP FN F1 $\uparrow$ Acc $\uparrow$
Stacking 20 Ours 29 59 10 2 0.84 0.88
Ensemble-DAgger 27 33 37 6 0.55 0.60
Thrifty-DAgger 33 30 34 3 0.49 0.63
Pushing (1 Expert) 50 Ours 29 51 9 11 0.74 0.80
Ensemble-DAgger 38 10 52 0 0.59 0.48
Thrifty-DAgger 37 6 57 0 0.57 0.43
Pushing (2 Experts) 25/25 Ours 27 51 8 14 0.71 0.78
Ensemble-DAgger 15 45 5 35 0.43 0.60
Thrifty-DAgger 25 38 15 22 0.57 0.63
Plugging 100 Ours 31 49 13 7 0.76 0.80
Ensemble-DAgger 22 37 18 23 0.52 0.59
Thrifty-DAgger 30 34 25 11 0.62 0.64
TABLE II: Confusion matrix comparing the binary robot query decisions (query or no query) with the binary task outcomes (success or failure) in 100 policy rollouts for visuomotor tasks. Here, True Positive (TP) is a query made to a failing episode. True Negative (TN) is a successful episode without query. False Positive (FP) is an unnecessary query made to a successful episode, and False Negative (FN) is a failure episode without query. ND is the number of expert demonstrations used for training the policy. The numbers are chosen to ensure the policy can both succeed and fail during task execution.
IV-A4 Choices to Accelerate Learning
Human-in-the-loop systems depend on the rapid training of robot policies, which necessitates fast policy learning. However, the original implementation of the diffusion policy is data-intensive, requiring hours of training [3]. In this section, we explore specific modifications that can significantly reduce the training time for diffusion policies.
First, the convergence rate of policy performance varies depending on the prediction type used for the diffusion model. We evaluate three prediction types—sample, epsilon, and V-objective—and find the V-objective [18] to be the best for training convergence. Secondly, using a small number of diffusion steps during training identical to the diffusion steps used for inference improves the convergence rate while not significantly affecting the policy performance. These choices enable the training to converge reasonably well in 300 epochs. Additionally, we set the observation horizon to one to reduce the amount of data processed in each batch. For image-based training, we use the lightweight R3M ResNet-18 model [26] and optionally freeze it to accelerate training at the cost of policy performance.
@@ IV-B Predictability of Robot Query on Task Failure
The effectiveness of a robot query system relies on how well it aligns with the actual need for expert intervention; Accurate predictions allow queries to be made only when necessary, avoiding both over-querying and under-querying. To quantitatively evaluate this, we use a confusion matrix between robot query decision and task outcome. For each task, we train the robot policy on a fixed number of expert demonstrations and rollout from 100 different initial configurations. We record task completion without expert intervention and whether the policy estimates high uncertainty during each episode, as described in Section III-C. All methods use the same set of thresholds, including $\alpha$ and K, as shown in Table IV.
The results show that Diff-DAgger achieves a significant improvement over Ensemble-DAgger and Thrifty-DAgger in terms of its ability to predict task failure: Compared to Ensemble-DAgger, Diff-DAgger achieves 47% higher F1 and 45% greater accuracy scores on average across the visuomotor tasks. Compared to Thrifty-DAgger, Diff-DAgger improves F1 scores by 37% and accuracy by 43%. Notably, when learning from two pushing experts, both Ensemble-DAgger and Thrifty-DAgger yield high rates of false negative: This is because the action variance present in the training dataset is high due to diverging plans (clockwise paths and counter-clockwise paths) given the same state, resulting in the policies’ insensitivity toward true OOD data. In contrast, Diff-DAgger allow significantly lower rates of false positive in other tasks, indicating its ability to handle multi-modality in demonstration data. The detailed results across tasks are shown in Table II.
Stacking Pushing (1 Expert) Pushing (2 Experts) Plugging
State Image State Image State Image State Image
Behavior Cloning 0.83 0.85 0.92 0.87 0.91 0.91 0.73 0.68
Ensemble-DAgger 0.90 0.87 0.90 0.80 0.66 0.58 0.57 0.53
Thrifty-DAgger 0.88 0.87 0.85 0.78 0.77 0.73 0.69 0.60
Ours 0.98 1.00 0.97 0.99 0.98 0.97 0.87 0.83
@@
TABLE III: Performance comparison across different tasks with state-based and image-based observations. We report the success rates using 100 environment configurations after the final iteration of DAgger. For each task, the number of expert demonstrations is fixed (detailed in Table IV) across methods. Our results show that Diff-DAgger consistently outperforms the other DAgger baseline methods given the same number of demonstrations from the expert.
Task Action To Tp Ta Nc RN18 PredTg Td Ni Nf Nd $\alpha$ K Nb
Stacking EEF 1 24 8 3 PT V-obj 8 20 52 4 0.99 2 256
Pushing Joint 1 24 2 1 FT Epsilon 8 40 92 4 0.98 1 256
Plugging EEF 1 48 16 3 FT V-obj 64 20 100 8 0.99 3 256
@@
TABLE IV: Hyperparameters for Diffusion Policy and DAgger. Action: action space, To: observation horizon, Tp: prediction horizon, Ta: rollout horizon, Nc: number of cameras, RN18: training strategy for vision encoder where PT means pre-trained and FT means fine-tuned, PredTg: prediction target when training diffusion policy, Td: number of diffusion timesteps for training and inference, Ni: number of initial expert demonstrations, Nf: final number of expert demonstrations, Nd: number of expert intervention after each training session, K: patience for consecutive violation of threshold, Nb: Diff-DAgger specific batch size for calculating the expected diffusion loss.

### v4

@@ Abstract
Recently, diffusion policy has shown impressive results in handling multi-modal tasks in robotic manipulation. However, it has fundamental limitations in out-of-distribution failures that persist due to compounding errors and its limited capability to extrapolate. One way to address these limitations is robot-gated DAgger, an interactive imitation learning with a robot query system to actively seek expert help during policy rollout. While robot-gated DAgger has high potential for learning at scale, existing methods like Ensemble-DAgger struggle with highly expressive policies: They often misinterpret policy disagreements as uncertainty at multi-modal decision points. To address this problem, we introduce Diff-DAgger, an efficient robot-gated DAgger algorithm that leverages the training objective of diffusion policy. We evaluate Diff-DAgger across different robot tasks including stacking, pushing, and plugging, and show that Diff-DAgger improves the task failure prediction by 39.0%, the task completion rate by 20.6%, and reduces the wall-clock time by a factor of 7.8. We hope that this work opens up a path for efficiently incorporating expressive yet data-hungry policies into interactive robot learning settings. The project website is available at: https://diffdagger.github.io.
@@ I Introduction
However, none of the existing robot-gated DAgger approaches fully addresses the multi-modality present in demonstration data. For example, the most widely adopted Ensemble-DAgger approaches use the action variance among an ensemble of policies to estimate uncertainty [6, 7, 8]. This means that when tasks admit multiple solutions, policies might erroneously indicate high uncertainty even when in a previously visited state. We illustrate this problem with a 2D navigation example (Figure 2 (a)), where two primary paths—clockwise and counter-clockwise around a circle—lead to the goal. This scenario allows us to categorize regions as in-distribution uni-modal (along the circular paths, shown as mixed red and orange), in-distribution multi-modal (at the path divergence point, shown as green), and out-of-distribution (significant deviations from expected paths, shown as blue). While ensemble methods for robot-gated DAgger effectively distinguish in-distribution uni-modal from out-of-distribution data using action variance, they fail at in-distribution multi-modal points due to policy disagreement (Figure 2 (b)).
@@
Figure 2: (a) Visualization of 2D navigation toy example. The orange dots represent the mode for following a clockwise path around the circle and the red dots represent another mode for following a counter-clockwise path. The green dots represent the multi-modal region where multiple diverging paths exist. The blue region indicates out-of-distribution. (b) Comparison of action standard deviation and diffusion loss for different regions across 100 runs in the 2D navigation example. The left bar plot displays the standard deviation of action across policies for in-distribution uni-modal (ID, UM), in-distribution multi-modal (ID, MM), and out-of-distribution (OOD) regions. The right bar plot shows the corresponding diffusion loss for these regions.
Incorporating an expressive policy class that handles multi-modal data into the DAgger framework is a promising but under-explored direction. In light of this, we propose Diff-DAgger, an efficient robot-gated DAgger algorithm that addresses multi-modality in demonstration data while reducing wall-clock time compared to other methods. Diff-DAgger uses a query system based on the loss function of the diffusion model, referred to as diffusion loss, which significantly improves the effectiveness of robot queries across various evaluation metrics. In five challenging manipulation tasks, including two real-world tasks, Diff-DAgger achieves a 39.0% higher F1 score in predicting task failure, and the collected data leads to 20.6% improvements in policy performance while improving the wall clock time by a factor of 7.8 in comparison to the baseline methods.
We make the following contributions: First, we develop a novel robot-gated query system based on diffusion loss to better handle multi-modality in expert demonstrations. Second, we quantitatively evaluate Diff-DAgger’s efficiency in terms of task failure prediction, task completion, and wall-clock time during interactive learning to demonstrate its superiority over other baseline methods. Lastly, we discuss design choices that accelerate policy learning to enable faster convergence and improved overall performance.
@@ II-B Uncertainty Estimation of Robot Action
The most widely used method for estimating action uncertainty is to use an ensemble of policies bootstrapped on a subset of the training data and using the level of disagreement among the policies output to estimate uncertainty [6, 7, 8]. However, the ensemble method for uncertainty estimation can be problematic in practice, especially when the task allows multiple modes of completion. Specifically, they are prone to either mistaking in-distribution multi-modal data as out-of-distribution data or vice versa. This is because expressive ensemble policies may disagree due to the presence of multiple viable strategies where each policy is certain about its generated plan, leading to high action variance. Furthermore, if the policy generates an action sequence instead of a single action, the uncertainty estimation may explode due to divergence over time or lose accuracy if a shorter horizon is used for calculating action variance.
@@
In contrast to the existing robot-gated DAgger methods, our approach handles complex data distribution by incorporating the loss function of diffusion policy into the decision rule. This approach does not need multiple policies for uncertainty estimation, reducing both training and inference time while benefiting from using an expressive policy class.
@@ III-B1 Action Diffusion
Diffusion Policy formulates action generation as a DDPM [2], where Langevin dynamics iteratively denoises a Gaussian prior, $\mathcal{N}(0,I)$, to recover structured actions. This approach enables stable training, handles multimodal action distributions, and has achieved state-of-the-art results in behavior cloning for robot manipulation [3].
@@ III-B2 Loss Function
During training, Diffusion Policy samples a denoising timestep and a Gaussian noise vector to corrupt the action sequence. Given the corrupted actions and the original observation sequence, it predicts the target objective and optimizes using the following loss function:
$L_{\pi}(o,a,\epsilon,t)=\|f(\epsilon,a,t)-f_{\theta}(o,\sqrt{\alpha_{t}}a+\sqrt{1-\alpha_{t}}\epsilon,t)\|^{2}$ (1)
where $o$ is the observation sequence, $a$ is the action sequence, and $\epsilon$ is a noise vector from Gaussian distribution $\mathcal{N}(0,I)$. The parameter $\alpha_{t}$ is a time-dependent factor that modulates the influence of the original data and the added noise during the diffusion process, and $t$ is the denoising timestep. The loss function $L(o,a,\epsilon,t)$ calculates the mean squared error between the prediction target $f(\epsilon,a,t)$, which depends on the choice of prediction type, and the noise predicted by the model $f_{\theta}$, given the noised data at time $t$ [18].
@@ III-C1 OOD Detection via Diffusion Loss
15:         if CDF(loss) $>\alpha$ for last $K$ timesteps then
@@ III-C2 Data Aggregation and Policy Update
We now summarize the data aggregation and policy update procedures. Diff-DAgger first initializes the policy $\pi_{r}$ by training it on Ni initial expert demonstrations. During rollout, the robot takes the actions in the environment until a high diffusion loss is detected. When the robot query is made, experts intervene from the current scene and complete the task. Upon completion of the task by either the robot or the expert, the task is initialized to a random initial configuration. If the robot neither queries nor completes the task after a time limit, automatic expert intervention takes place. After every Nd expert intervention has been made, we perform a policy update, which consists of policy training and threshold setting. The policy training involves training the policy for a fixed number of epochs using the previously collected expert demonstration data, and threshold setting involves setting the base values by iterating through the training dataset to set the OOD violation threshold used for robot-query trigger. The full pseudo-code is available in Algorithm 1.
@@ IV-A Experiment Setup
Figure 3: Overview of simulation and real-world tasks: stacking, pushing, plugging, and loading.
Type Task # Obj NC Contact Rich High Prec TL Expert Type Task Dur Succ Rate
Sim Stacking 2 3 False False 250 Motion Planner 60.3 1.00
Pushing 1 1 True False 250 RL Agent 65.7 0.93
Plugging 2 3 False True 500 Motion Planner 147 0.98
Real Stacking 2 3 False False 600 Human 174.2 1.00
Loading 2 3 True True 600 Human 210.5 1.00
TABLE I: Task and Expert Summary. # Obj: number of objects, NC: number of cameras, Contact Rich: contact-richness, High Prec: high-precision, TL: time limit in timesteps at 20 Hz in simulation or 30 Hz in real world, Task Dur: average expert task duration in simulation steps, Exp SR: expert success rate during intervention.
@@ IV-A1 Tasks and Experts
We test Diff-DAgger in the simulated and real robot environments. We include three tasks in simulation—stacking, pushing, and plugging—using the ManiSkill environment [19] and two tasks—stacking and loading—in real world. In the stacking task, the robot picks up a red cube and places it on a green cube, evaluating its basic perception and manipulation skills. The pushing task requires using an end-effector stick to move a red tee to a predefined position, requiring multimodal planning and control of contact dynamics. The plugging task involves inserting a charger into a receptacle, testing fine manipulation and alignment accuracy. Lastly, the loading task requires the robot to pick up the soft bread and precisely load in onto a toaster, assessing the robot’s ability to handle deformable objects with precision.
For simulation tasks, we design experts as either multi-modal or long-horizon planners to evaluate different expert types. Specifically, we use RL agents and motion planners as the experts: the motion planner, is a long-horizon planner that generates actions based on future goals, whereas multiple RL agents with different reward functions provide different modes for task completion. For stacking and plugging tasks, we use a rule-based RRT motion planner to simulate the expert [20]. For pushing, two experts are trained using PPO [21] with varying rewards: in addition to task-related reward, one agent is additionally rewarded for rotating the object clockwise direction, whereas other agent is rewarded for rotating counter-clockwise. During interactive learning, the agents alternate in providing demonstrations. We additionally compare learning from the counter-clockwise agent only. For real-world tasks, a proficient human expert teleoperated the robot using a VR controller. The tasks and experts are summarized in Table I.
For the task setup, we use three cameras—left view, right view, and in-hand view—for pick-and-place tasks to mitigate visual occlusion, while a single top-down camera is used for the pushing task. In real-world experiments, we use a Franka Research 3 robot and use three RealSense D435 cameras, capturing 720p RGB images at 30 FPS, which is then downsampled to 256p at 15 FPS.
@@ IV-A2 Baselines
IV-A3 Model Learning
We adopt a model architecture similar to [22], utilizing a U-Net [22] policy head with FiLM conditioning [23]. We test both state-based and image-based data: For the state-based case, we use the internal robot proprioception data (robot joint position and end-effector pose) and the ground-truth pose of the external objects as the input. For the image-based case, we apply random crop augmentation following [24] and extract the embedding vector using an R3M ResNet18 encoder [25]. Instead of average pooling, we apply spatial softmax [26] to the extracted features before concatenating them with proprioception data. Additionally, the observation horizon is set to one to reduce the amount of data processed in each batch. All approaches use position-based control, using either joint position control or end-effector pose control, with either absolute or relative action representations [27]. For end-effector pose control, a 6D representation for the orientation is used [28].
@@ IV-B Predictability of Robot Query on Task Failure
The effectiveness of a robot query system depends on its ability to respond to a failing episode. To quantitatively assess this, we evaluate it in simulation, where task outcomes can be efficiently recorded without the need for real-world rollouts. Specifically, we construct a confusion matrix between the robot’s query decisions and task outcomes: For each task, we train the robot policy on a fixed number of expert demonstrations to roughly match 50% success rate and evaluate for 100 randomly initialized episodes. We record task completion without expert intervention and whether the policy estimates high uncertainty during each episode, as described in Section III-C. For each task, all methods use the same set of thresholds, including $\alpha$ and K, as shown in Table IV.
The results show that Diff-DAgger achieves a significant improvement over Ensemble-DAgger and Thrifty-DAgger in terms of its ability to predict task failure: Compared to Ensemble-DAgger, Diff-DAgger achieves 42.7% higher F1 and 38.4% greater accuracy scores on average across the simulated visuomotor tasks. Compared to Thrifty-DAgger, Diff-DAgger improves F1 scores by 22.6% and accuracy by 32.8%. The results across tasks are shown in Table IV.
Task ND Method TPR TNR F1 $\uparrow$ Acc $\uparrow$
Stacking 30 Ours 0.88 0.77 0.83 0.84
Ensemble-DA 0.81 0.31 0.52 0.59
Thrifty-DA 0.95 0.21 0.54 0.65
Pushing (1 Expert) 40 Ours 0.83 0.92 0.88 0.85
Ensemble-DA 1.00 0.11 0.60 0.73
Thrifty-DA 1.00 0.02 0.57 0.72
Pushing (2 Experts) 25/25 Ours 1.00 0.52 0.77 0.82
Ensemble-DA 0.29 0.90 0.58 0.42
Thrifty-DA 0.58 0.66 0.62 0.62
Plugging 100 Ours 0.68 0.82 0.75 0.73
Ensemble-DA 0.22 0.27 0.14 0.37
Thrifty-DA 0.26 0.31 0.22 0.21
TABLE II: Comparison of robot query decisions (query or no query) against task outcomes (success or failure) in 100 policy rollouts for visuomotor tasks. Here, True Positive Rate (TPR) refers to the proportion of queries made to failing episodes, while True Negative Rate (TNR) refers to the proportion of non-queries made to successful episodes. ND is the number of expert demonstrations used for training the policy.
Stacking (Sim) Pushing (1 Expert) Pushing (2 Experts) Plugging Stacking (Real) Loading
State Image State Image State Image State Image Image Image
BC 0.89 0.85 0.89 0.72 0.74 0.71 0.75 0.49 0.6 0.8
Ensemble-DA 0.72 0.83 0.87 0.81 0.70 0.59 0.69 0.37 - -
Thrifty-DA 0.79 0.86 0.85 0.81 0.69 0.68 0.65 0.47 - -
Ours 0.94 0.94 0.96 0.87 0.87 0.79 0.92 0.84 0.8 0.95
@@
TABLE III: Performance comparison across different tasks with state-based and image-based observations in simulation or real-world. We report success rates after the final DAgger iteration, using 100 evaluation episodes for simulation tasks and 20 for real-world tasks. For each task, the number of expert demonstrations is fixed (detailed in Table IV) across methods. Our results show that Diff-DAgger consistently outperforms the other methods given the same number of demonstrations from the expert.
Task Action To Tp Ta PredTg Td Ni Nf Nd $\alpha$ K Nb
Stacking (Sim) Abs Pose 1 32 8 V-obj 16 20 60 4 0.99 2 512
Pushing (1 Expert) Rel Joint 1 32 2 V-obj 16 20 100 4 0.99 1 512
Pushing (2 Experts) Rel Joint 1 32 2 V-obj 16 20 100 4 0.9 2 512
Plugging Abs Pose 1 64 8 V-obj 64 20 100 8 0.99 2 512
Stacking (Real) Abs Pose 1 32 8 V-obj 16 40 120 20 0.99 2 512
Loading Abs Pose 1 64 8 V-obj 64 40 120 20 0.99 2 512
@@
TABLE IV: Hyperparameters for Diffusion Policy and DAgger. Action: action space, To: observation horizon, Tp: prediction horizon, Ta: rollout horizon, PredTg: prediction target when training diffusion policy, Td: number of diffusion timesteps for training and inference, Ni: number of initial expert demonstrations, Nf: final number of expert demonstrations, Nd: number of expert intervention after each training session, K: patience for violation, Nb: Diff-DAgger specific batch size for calculating the diffusion loss.